OpenStudy (anonymous):

Solve each quadratic equation by factoring 2x^2 + 4x = 6 x^2 - 6x = 16 2x^2 + 4x - 70 = 0 2x^2 - 72 = 0

OpenStudy (anonymous):

Like, 2(x^2+2x+1)-8=0, 2(x+1)^2-8=0, (x+1)^2-4=0, (x+1)^2-2^2=0, (x+1+2)(x+1-2)=0, (x+3)(x-1)=0, x=-3 or x=1?

jhonyy9 (jhonyy9):

x2 -6x =16 x2 -6x -16=0 (x-8)(x+2)=0

jhonyy9 (jhonyy9):

so than x_1 = 8 and x_2 = -2

jhonyy9 (jhonyy9):

2x2 +4x-70 =0 divide by 2 x2 +2x -35=0 x_1,_2=(-2+/- sqrt(4+140))/2=(-2+/- 12)/2=-7 and 5 factorized is (x+7)(x-5) so x_1,_2= -7 and 5

jhonyy9 (jhonyy9):

2x2 -72=0 divide by 2 x2 -36=0 factorized is (x-6)(x+6) so x_1,_2 = -6 and 6

jhonyy9 (jhonyy9):

the first is 2x2 +4x =6 2x2 +4x -6=0 divide by 2 x2 +2x -3=0 factorized is (x+3)(x-1) so x_1,_2= -3 and 1